# Enhanced Single-Photon Detector: Achieving Superconducting-Level Performance with Conventional Quantum Technology

### v1

@@ 3.1 Device Requirements
Finally, the requirements for the initial single-photon detector (SPD) used for initialization are minimal. As demonstrated in Section 2.4, the stable performance of the ESPD is largely independent of the specific characteristics of the starting detector, provided it falls within a certain convergence interval. The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point. Moreover, while a trade-off exists between DE and DCR, the DCR of the ESPD can be drastically reduced by increasing $k$, and the DE can be significantly enhanced by selecting a small $k$. One or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient. Indeed, simulations confirm that both an SPD with low DE and low DCR and one with moderate DE and high DCR serve as viable starting points. In summary, the ESPD scheme only requires a normal, low-performance SPD for initialization, which is easily achieved. For instance, a device dating back two decades [28] is sufficient.
3.2 Applications
To illustrate the practical impact of the ESPD scheme, we consider its application in QKD. For a QKD protocol having a secure threshold $e_{th}$, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by:
Given that the ESPD paradigm reduces the DCR ($d$) by several orders of magnitude, it proportionately lowers the minimal tolerable channel transmission rate $\gamma$ by the same factor. Furthermore, the simultaneous enhancement of the DE ($\eta$) not only lowers $\gamma$ but also translates directly into an improved QKD key rate at any fixed distance.
Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD scheme avoids introducing an extra sifting term that penalizes efficiency. This allows for significant QKD performance improvements without an additional efficiency cost. Consequently, the ESPD paradigm offers a compelling first-choice solution for practical quantum communication implementations.
3.3 Limitation and Future Work
@@
Despite the significance of the ESPD framework, some limitations remain that are left for future investigations.
@@ 3.3.1 Theoretical Completeness
The theoretical framework established in Section 2.2 remains an area for deeper exploration. While the numerical designs presented in Section 2.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest. Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required. Given the inherent mathematical difficulty of such a non-linear and time-variant system, these specialized studies are reserved for the future.
3.3.2 Practical Cost and Optimization
Although the parallelized nature of the ESPD’s detection stages theoretically preserves detection efficiency, the requirement for multiple detection events introduces a practical cost and operational burden, even without increasing the fundamental technological barriers. An $L$-level ESPD, with $n$ auxiliary signals per level, requires a total of $(n+1)^{L}$ base detections. While effective performance is typically achieved with only a few levels, this inherent multiplicative complexity increases the practical cost, which may limit immediate commercial scalability. As shown in Section 2.4, the optimal choice of $n$ and $k$ is crucial as it dictates the convergence speed and the resulting DE-DCR trade-off. Enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and larger $n$, and enlarging $n$ increases detection cost. Therefore, designing the ESPD with optimal parameters constitutes a significant multi-objective optimization problem for the future.
@@ 3.3.2 Practical Cost and Optimization
This optimization problem also encompasses the optimal circuit design. To mitigate error propagation, minimizing the maximal circuit length is essential. Optimal designs could involve leveraging a binary-tree structure and defining the positive report event via a global, rather than the current level-by-level local, decision criterion, which expands the tree structure used in Section 2.1 with a local decision criterion. However, both problems require further and specific discussions: the general binary-tree structure results in non-independent distributions that are difficult to analyze, while the global decision criterion design must account for both device requirements and implementation convenience.
@@ 3.3.3 Physical Integrations and Experiments
Despite the convincing demonstration of effectiveness via numerical simulations in Section 2.4, the physical implementation of the ESPD scheme requires dedicated experimental validation. Furthermore, the seamless integration of controlled gates into various quantum optical tasks has not yet been fully investigated. Thus, both the physical realization of the ESPD scheme and its application to a wide range of tasks are key directions for future work. Reassuringly, as established in Section 3.1, no fundamental technological barriers prevent its current achievement.
@@ 4 Conclusion
In this paper, we have introduced the Enhanced Single-Photon Detector (ESPD) framework, a robust methodology for transforming low-performance, conventional SPDs into devices with performance comparable to or exceeding state-of-the-art superconducting SPDs. The ESPD scheme is characterized by significantly reduced technology requirements and immediate physical realizability using current quantum components. Our numerical analysis demonstrates that a legacy non-superconducting SPD, with DE $\approx 59\%$ and DCR $\approx 10^{-2}$, can be enhanced to one achieving DE $95\%$ and DCR $10^{-9}$. This dramatic boost circumvents the need for cryogenic superconducting materials, enabling the deployment of high-performance detection capabilities outside specialized laboratories. Furthermore, this enhancement allows the minimal tolerable channel transmission rate in quantum communication to be reduced by multiple orders of magnitude. The ESPD framework thus paves a clear, scalable path toward the commercialization and widespread implementation of high-performance quantum optical tasks.
@@
The author declares no financial or non-financial competing interests.

### v2

@@ 3.1.3 Choice of initialized SPDs
It is also important to note that the requirements for SPD used for initialization are minimal, further lowering the barrier to implementation. As demonstrated in comparing Table 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the starting detector, provided it falls within a certain convergence interval. The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point. Moreover, while a trade-off exists between DE and DCR, the DCR of the ESPD can be drastically reduced by increasing $k$, and the DE can be significantly enhanced by selecting a small $k$. Therefore, one or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient. Indeed, simulations confirm that both an SPD with low DE and low DCR and one with moderate DE and high DCR serve as viable initialization.
In summary, the ESPD scheme only requires a normal, room-temperature SPD for initialization, which is routinely achievable. For instance, even an SPD developed two decades ago [28] is sufficient.
3.1.4 Auxiliary Sources
Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, we note that implementing the ESPD scheme may not require single photon sources. The key parameter $P$ represents the detection probability of the state $|1\rangle$ from the auxiliary path with an ideal SPD when the control signal input is non-vacuum, regardless of the exact number of photons or the presence of unwanted $|0\rangle$. Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more opportunities for detecting the required $|1\rangle$ [^3] Since the employment of the filter block, for a non-vacuum input of a controlled module, if the auxiliary multi-photon signal results in more state $|1\rangle$, then a positive report has a higher chance to be produced, which provides benefits since it enlarges $P$, while, if the auxiliary multi-photon signal adds state $|0\rangle$, then they will be filtered out, thus providing no effects. Similarly, for a vacuum input of a controlled module, only $|0\rangle$ is added in auxiliary sources, which will be filtered out and thereby have no effect even if the source is multiple-photon.. However, the formulation of this, including its impact on controlled operations, needs a rigorous discussion and is reserved for future work.
3.2 Limitations and Future Works
@@
Despite the significance of the ESPD framework, some limitations remain, which are left for future investigations.
@@ 3.2.1 Theoretical Completeness
The theoretical framework established in Section 2.2 remains an area for deeper exploration. While the numerical designs presented in Section 2.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest. Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required. Given the inherent mathematical difficulty of such a non-linear and time-variant system and the sufficiency for practical technology employments by numerical designs, these specialized mathematical studies are reserved for the future.
3.2.2 Resource Overhead vs. Environmental Simplicity
While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels. Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections. Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD. For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (if not applying techniques such as multiplexing), resulting in hundreds of integrated components. While not posing a fundamental physical barrier (as discussed in Section 3.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
However, this complexity is well within the capabilities of modern large-scale photonic integrated circuits. Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57]. Notably, very-large-scale integration of quantum photonics has been experimentally realized with over 2,500 components on a single silicon-on-insulator chip [58], verifying that the component density required for a multi-level ESPD is manufacturable using standard complementary metal-oxide-semiconductor processes. Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss material platforms. For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
In summary, the hardware complexity of the ESPD should be evaluated as a deliberate engineering trade-off against the necessity of cryogenic operation. The SNSPD paradigm mandates bulky, energy-intensive, and expensive cryogenic infrastructure (typically operating below 4 K). The ESPD scheme, conversely, shifts the burden from a demanding operating environment to the domain of on-chip integration. By leveraging large-scale photonic platforms, thousands of components can be lithographically fabricated on a single chip, effectively converting high operational costs and environmental constraints into a one-time design and fabrication cost. This trade-off is particularly advantageous for applications where size, weight, and power constraints prohibit the use of cryogenics, such as satellite-based quantum communication payloads, mobile quantum nodes, or research laboratories lacking dedicated low-temperature facilities.
3.2.3 Optimization Design
The inherent exponential resource scaling discussed in Section 3.2.2 necessitates optimization for future practical deployments. Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$). Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence. As shown in Section 2.4, the optimal choice of parameters $n$ (number of auxiliary signals in each level) and $k$ (decision threshold) is crucial: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost. Furthermore, as shown in Table 7 and visualized in Fig.7, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
Figure 7: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different $k$. Para 3 is the same as in Fig.2.
Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section 2.4
Para 4 Para 5 Para 6
[…]
1 (8,2) $(95.0\%,\ 4.2\times 10^{-3})$ (4,2) $(86.1\%,\ 1.2\times 10^{-3})$ (3,1) $(95.8\%,\ 4.3\times 10^{-2})$
2 (8,4) $(93.5\%,\ 1.4\times 10^{-7})$ (8,4) $(92.6\%,\ 6.4\times 10^{-9})$ (6,4) $(93.9\%,\ 1.2\times 10^{-4})$
3 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.3\%,\ 8.2\times 10^{-10})$ (8,4) $(93.4\%,\ 1.2\times 10^{-9})$
4 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.3\%,\ 8.4\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
5 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
6 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
7 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
8 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
@@ 3.2.3 Optimization Design
The optimization problem also encompasses the optimal circuit topology. To mitigate cumulative error propagation, minimizing the maximal circuit length is essential. As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the design: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section 2.1. However, these architectural changes introduce new theoretical challenges: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
@@ 3.2.4 Physical Integrations and Experiments
Despite the convincing demonstration of effectiveness via numerical simulations in Section 2.4, physical implementations of the ESPD scheme require dedicated experimental validations. Reassuringly, as established in Section 3.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization. However, the realization of the large-scale iterative circuit demands significant and focused engineering efforts, particularly concerning the precise integration and synchronization of the cascaded components. Furthermore, the seamless integration of the enhanced detector into various quantum optical tasks (such as QKD) needs full investigation. Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
3.2.5 Count Rate and Timing Jitter
While the ESPD framework significantly enhances DE and suppresses DCR, it generally inherits the timing characteristics of the baseline detectors. Specifically, the system’s timing jitter and maximum count rate remain constrained by the intrinsic performance of the initial SPDs (e.g., the avalanche buildup time and dead time of InGaAs/InP devices). However, the architecture does not impose significant additional penalties: the optical-controlled quantum operations function on timescales usually faster than the electronic response of room-temperature SPDs, ensuring negligible added temporal broadening. Furthermore, as the final SPDs operate in parallel, the effective system throughput remains comparable to that of the constituent devices. Nevertheless, future iterations of the ESPD scheme could incorporate techniques such as multiplexing strategies to optimize these temporal metrics.
@@ 4 Conclusion
We have introduced the Enhanced Single-Photon Detector (ESPD) framework, a robust methodology for upgrading low-performance SPDs into devices whose capabilities can match or even surpass those of state-of-the-art superconducting SPDs. The distinguished advantage of the ESPD paradigm is its foundation on non-superconducting components entirely, obviating the demanding cryogenic operational and infrastructural requirements of SNSPDs, and thereby enabling high-performance SPDs to be utilized beyond specialized low-temperature laboratories. Our numerical analysis shows that a legacy non-superconducting SPD (with $\text{DE}\approx 59\%$ and $\text{DCR}\approx 10^{-2}$) can be enhanced to achieve $\text{DE}>93\%$ and $\text{DCR}<10^{-9}$, reducing the minimal tolerable channel transmission rate in quantum communication by several orders of magnitude. Although physically realizing such a large-scale architecture requires substantial integration and engineering efforts, the ESPD framework represents a paradigm shift, trading resource overhead for operation-environmental simplicity. It opens the door to high-performance quantum-optical applications without cryogenic cooling, providing a clear and viable path toward the widespread implementation of quantum optical tasks in condition-constrained environments.
@@
The author declares no competing interests.
Funding Statements
There is no funding for this work.